OpenStudy (anonymous):

Use substitution to solve the system 3x − y = 5 2x + 3y = −4 (x,y)= ( , )

OpenStudy (anonymous):

solve the first eqn for y y=3x-5 (*) sub this expression for x into eqn 2 2(3x-5)+3y=-4 solve for y 6x-10+3y=-4 9x=6 x=2/3 sub 2/3 for x in (*) y=3(2/3)-5=2-5=-3 the solution is (2/3, -3)

OpenStudy (anonymous):

lol i did a problem ezactly like tis on a quiz today...

OpenStudy (phi):

y=3x-5 (*) should go into 2x + 3y= -4

OpenStudy (phi):

so you should solve 2x + 3(3x-5) = -4

OpenStudy (anonymous):

ur correct; i think its time to quite for the night!

OpenStudy (phi):

It's easy to get wrapped around the axle sometimes!

OpenStudy (anonymous):

good night mandolino

OpenStudy (anonymous):

solve the first eqn for y y=3x-5 (*) sub this expression for x into eqn 2 2x+3(3x-5)=-4 solve for y 2x+9x-15=-4 11x=11 x=1 sub 1 for x in (*) y=3(1)-5=3-5=-2 the solution is (1, -2)

OpenStudy (anonymous):

g'night :})

OpenStudy (anonymous):

thanks guys!!!

OpenStudy (anonymous):

and goodman what school do u go to??

OpenStudy (anonymous):

JAMES LOGAN HIGH SCHOOL

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

i was thinking we went to the same college because everyone has to take the same quize over the internet in my class, that has the same questions and its due tonight at like midnight, i just thouht there could be a chance, that would be sweet if true

OpenStudy (anonymous):

lol..college..im only a freshman still good to know u!

OpenStudy (anonymous):

you could easily be in my college class, i bet u would think it easy, its math 110, but its hard for me,my tallents not math

OpenStudy (anonymous):

me too..im in alg 2 and have the worst teacher...